# Box Confidence Depth: simulation-based inference with hyper-rectangles

### v1

@@ 3.5 High dimensional hyper rectangles
Denote as ${B}_{t}^{*(d)}$ a random Box based on dimension $j\in\{1,\ldots,d\}$. Then
$Pr(t^{\text{obs}}\in\mathcal{B}_{t}^{*(d-1)})\leq Pr(t^{\text{obs}}\in\mathcal{B}_{t}^{*(d)}).$
@@
The previous statement means that the acceptance probability reduces when the dimension of the statistic grows. This leads to challenges in accurately estimating the tails of the Box-CD function, as the corresponding parameter regions are associated with rare events, leading to high variance of the procedure.
@@
To alleviate the problem, without compromising the validity of the procedure, we introduce a generalization of the Box-CD approach based on generating a series of $S$ pseudo-samples instead of a pair. We only require that $S$ is a even number. Define
@@
The computational cost increases by $S/2$ while accepted values might grow faster than this rate, enabling better estimation of the function, especially in the tails. In Section 4, through empirical analysis, we demonstrate that adopting this second strategy is preferable when the dimension of the summary statistic increases to generating a larger number of candidate values solely.
@@ 4 Examples
where $h>0$ serves as the bandwidth parameter, internally chosen via cross-validation, and $K(\mathbf{u})=\prod_{j=1}^{d}k(u_{j})$, with $k(u_{j})$ representing the univariate Gaussian kernel function independently applied across each dimension $j$. We employed a 1-Nearest neighbor method to assess whether $\theta_{0}$ was included in the confidence regions, by predicting the value of $\mathcal{CD}^{\text{box}}(\theta_{0})$.
@@ 4.3 Mixture
Consider the normal mixture model $y\sim 0.5\mathcal{N}(-\theta,1)+0.5\mathcal{N}(\theta,1).$ The summary statistic used is the ordered sample $t=(y^{\text{obs}}_{(1)},\ldots,y^{\text{obs}}_{(n)})$, of size $d=n$. The proposal for $\theta$ is a $\text{Uniform}[0,3]$ and we fix $n=10,15,20,25$. We use this model to study the acceptance ratio as a function of the dimension of the summary statistics ($d$) and the number of pseudo-samples, governed by the hyper-parameter $S$. Figure 3 reports the absolute number of accepted proposals from draws of size $R=100000$ (left) and ratio compared to accepted proposals with $S=2$. For a fixed $S$, the number of accepted draws decreases as $d$ increases, causing loss in efficiency. When $S$ varies, the number of accepted parameters is not proportional to $R$, but grows faster (see the right panel of Figure 3).
@@
Figure 4: Five replications of the same Box-CD function for the position parameter in the mixture model, with number of pseudo-samples $S$ varying. The vertical line indicates the true generating parameter $\theta_{0}=0.8$.
Figure 4 presents the Box-CD functions derived from a sample of size $n=25$ drawn using as a true generating parameter $\theta_{0}=0.8$. This study considers four different scenarios with pseudo-sample sizes of $S=4,6,8,10$. For every value of $S$, five replications of the Box-CD are generated using the same observed sample. When $S$ grows, the procedure’s variability diminishes. Note that the tails of the function become heavier as $S$ grows. This phenomenon poses no problem as demonstrated by the simulation study (Table 1) because the confidence sets are reliant on the value of the function $CD(\theta)$ instead of tail areas. In conducting the simulation study, we set $n=d=10$ and $S=6$.
@@ 4.4 Ricker’s Model
We conduct two experiments: first, we assume that only the log-growth rate is unknown and consider as summary statistics the median of counts and the quantiles of level 0.25 0.75. For the second experiment, both the parameters $log(r)$ and $\sigma^{2}$ were considered unknown, and we used as the set of summary statistics the whole time series minus the first observation, thus of length $d=19$. The number of pseudo-samples for each proposals were $S=2$ and $S=10$ in the two experiments, respectively. The empirical coverages of the confidence sets are conformal with the nominal (Table 1). Two examples of confidence regions obtained for two independent draws from the model with parameters $log(r)=2$ and $\sigma^{2}=2$ are reported in Figure 5.

### v2

@@ 3.6 High dimensional hyper rectangles
Denote as $\mathcal{B}_{t}^{*(d)}$ a random Box based on dimension $j\in\{1,\ldots,d\}$. Then
$Pr(t^{\text{obs}}_{1},\ldots,t^{\text{obs}}_{d-1}\in\mathcal{B}_{t}^{*(d-1)})\geq Pr(t^{\text{obs}}_{1},\ldots,t^{\text{obs}}_{d}\in\mathcal{B}_{t}^{*(d)}).$
This inequality reflects that the acceptance probability decreases as the dimensionality of the statistic increases. Intuitively, for the observed point $t^{\text{obs}}$ to lie within the $d$-dimensional box $\mathcal{B}_{t}^{*(d)}$, all $d$ components must fall within their respective coordinate-wise intervals. This implies that any subset of d1d-1 components must also fall within their corresponding intervals. In contrast, for the (d1)(d-1)-dimensional case, only a subset of these constraints needs to be satisfied, making it more likely for the point to be accepted.
@@
This leads to challenges in accurately estimating the tails of the Box-CD function, as the corresponding parameter regions are associated with rare events, especially in high dimensions, as it happens for ABC.
However, as stated in Section 3, a crucial property of distance and divergence functions typically used in ABC is the identity of indiscernible. This condition is not fulfilled by the type of discrepancy associated to the acceptance criterion of Box-CD. This would only occur in a degenerate case where the set $\mathcal{B}^{*}_{t}$ is a singleton with probability one, a scenario not endorsed by our assumptions. The failure to meet this property carries significant implications. In particular, it allows the method to accept parameter values even when the simulated summary statistics are far from the observed ones, as long as they fall within a coarse acceptance region. This does not rule out the problem of the curse of dimensionality. But we cannot study this issue straight under the lens of the distance-based approaches. To illustrate this, consider this simplified scenario. Let $t^{\text{obs}}\in\mathbb{R}^{d}$ be a fixed observed summary statistic, from the model $\mathcal{N}_{d}(0,I_{d})$ and let us assume that $t\sim\mathcal{N}_{d}(0,V)$ is a simulated statistic from the prior predictive distribution, where $V=v\cdot I_{d}$ is diagonal. In ABC, the common acceptance criterion is based on the fixed-radius ball
$\|t-t^{\text{obs}}\|\leq\epsilon,$
for some fixed tolerance $\epsilon>0$. Then the acceptance probability satisfies the identity property. Now consider without loss of generality $t^{\text{obs}}=0_{d}$ and $t$ independent random vectors from a standard multivariate normal distribution in $\mathbb{R}^{d}$. We are interested in the probability that their Euclidean distance is less than or equal to $\epsilon$, i.e.
$\mathbb{P}(\|t-t^{\text{obs}}\|\leq\epsilon).$
Let us denote with $\Delta=t_{1}-t_{2}$ the difference between $t_{1}$ and $t_{2}$. Since $t_{1}$ and $t_{2}$ are independent and both distributed as $\mathcal{N}_{d}(0,I_{d})$, the difference $\Delta$ is distributed as $\Delta\sim\mathcal{N}_{d}(0,V+I_{d})$. Therefore, the squared norm follows the scaled chi-squared distribution
$\|\Delta\|^{2}\sim(v+1)\cdot\chi^{2}_{d}.$
Hence, the probability of interest becomes
$\mathbb{P}(\|t-t^{\text{obs}}\|\leq\epsilon)=\mathbb{P}((v+1)\cdot\chi^{2}_{d}\leq\epsilon^{2})=F_{\chi^{2}_{d}}\left(\frac{\epsilon^{2}}{v+1}\right),$
where $F_{\chi^{2}_{d}}(\cdot)$ is the cumulative distribution function (CDF) of the chi-squared distribution with $d$ degrees of freedom. The expectation and variance of $\|\Delta\|^{2}$ are, respectively,
$\mathbb{E}(\|\Delta\|^{2})=(v+1)d\quad\text{and}\quad\text{Var}(\|\Delta\|^{2})=(v+1)^{2}d.$
So, as $d\to\infty$, the distance grows roughly as $\sqrt{(v+1)d}$. For fixed $\epsilon$, the probability decays rapidly with $d$. Moreover, the density of $\Delta\sim\mathcal{N}_{d}(0,(v+1)I_{d})$ is approximately constant close to zero (as we are interested in having $\epsilon\approx 0$). Denoting with $Vol(\epsilon)$ the volume of the ball with radius $\epsilon$, then
$\mathbb{P}(\|t-t^{\textbf{obs}}\|\leq\epsilon)\approx p(t^{\text{obs}})\cdot Vol(\epsilon)=p(t^{\text{obs}})\cdot\frac{\pi^{d/2}}{\Gamma(d/2+1)}\cdot\epsilon^{d},$
with $\Gamma\left(\frac{d}{2}+1\right)\approx\sqrt{\pi d}\left(\frac{d}{2e}\right)^{\frac{d}{2}}$ by Stirling’s approximation. Hence for large $d$
$\mathbb{P}\big(\|t_{1}-t_{2}\|\leq\epsilon\big)\propto\frac{1}{\sqrt{d}}\left(\epsilon\frac{\sqrt{2\pi e}}{\sqrt{d}}\right)^{d}<\epsilon^{d},$
which decays super-exponentially for the presence of $(1/\sqrt{d})^{d}$. Alternatively, consider the acceptance region defined by a data-adaptive axis-aligned hyperrectangle:
$\mathcal{B}(t^{(1)},t^{(2)})=\{t\in\mathbb{R}^{d}:\min(t^{(1)}_{j},t^{(2)}_{j})\leq t_{j}\leq\max(t^{(1)}_{j},t^{(2)}_{j})\ \text{for all }j=1,\ldots,d\},$
where $t^{(1)},t^{(2)}\sim\mathcal{N}_{d}(0,V)$ are two independent simulated summaries from the prior predictive. Then, the acceptance probability is
$\mathbb{P}(t^{\text{obs}}\in\mathcal{B}(t^{(1)},t^{(2)}))\propto\frac{1}{\prod_{j=1}^{d}(1-x_{j})x_{j}}\leq\left(\frac{1}{4}\right)^{d},$
with $0<x<1$ and $x=1/2$ if the prior/proposal is symmetric and centered around the the point of maximal depth. This quantity decays exponentially with the dimension, still providing a better rate than ABC.
@@
To alleviate the problem of low acceptance probability that translates into inefficiency in high dimensions, we introduce a generalization of the Box-CD approach based on generating a series of $S$ pseudo-samples instead of a pair, without compromising the validity of the procedure. We only require that $S$ to be a even number. Define
@@
The computational cost increases by $S/2$ due to the larger number of model simulations; however, the number of accepted values may grow faster than this rate, allowing for more accurate estimation of the target function—particularly in the tails—under a fixed computational budget. In particular, increasing $S$ offers a natural opportunity to exploit parallel computation: using $S$ parallel processors can in principle increase the probability of accepting beyond a factor of $S$. This contrasts with standard ABC rejection techniques, which typically cannot exploit multiple proposals as effectively. In ABC rejection, imposing a joint distance condition on multiple proposals alters the likelihood and decreases the acceptance probability; even when multiple model proposals are generated in parallel, and conditions are imposed independently, the number of accepted samples can only scale linearly with the number of proposals. In Example 4.3 we empirically examine the effect of choosing $S>2$, with particular attention to scenarios where the dimension of summary statistics increases.
@@ 4 Examples
where $h>0$ is the bandwidth parameter, internally chosen via cross-validation, and $K(\mathbf{u})=\prod_{j=1}^{d}k(u_{j})$, with $k(u_{j})$ representing the univariate Gaussian kernel function independently applied across each dimension $j$. We employed a 1-Nearest neighbor method to assess whether $\theta_{0}$ was included in the confidence regions, by predicting the value of $\mathcal{CD}^{\text{box}}(\theta_{0})$.
@@ 4.2 Multivariate $t$ distribution
In this example the summary statistics, in addition to being dependent, have a distribution that depends on all the unknown parameters. By contrast, if the scale parameters were unknown but the summaries included only the means, the acceptance probability would not reach the upper bound of $1$ for increasing scales, violating the assumptions. If the correlation parameters were unknown instead, by using only the empirical means as summaries, the acceptance probability would be constant marginally in the correlations - again violating the assumptions and leading to non-informative regions whose depth is constant.
@@ 4.3 Mixture
[…]
We focus on this model to study the acceptance ratio as a function of the dimension of the summary statistics ($d$) and the number of pseudo-samples, governed by the hyper-parameter $S$. Figure 3 reports the total number of accepted proposals from draws of size $R=100000$ (left) and ratio compared to accepted proposals with $S=2$. For a fixed $S$, the number of accepted draws decreases as $d$ increases, causing loss in efficiency. When $S$ varies, the number of accepted parameters is not proportional to $R$, but grows faster (see the right panel of Figure 3).
@@
Figure 4: Five replications of the same Box-CD function, with fixed $y^{\text{obs}}$ for the position parameter in the mixture model, with number of pseudo-samples $S$ varying. The vertical line indicates the true generating parameter $\theta_{0}=0.8$.
Figure 4 presents the Box-CD functions derived from a sample of size $n=25$ drawn using as a true generating parameter $\theta_{0}=0.8$. For every value of $S=4$ and $S=10$, five replications of the Box-CD are generated using the same observed sample. When $S$ grows, the procedure’s variability diminishes. Note that the tails of the function become heavier as $S$ grows. This phenomenon poses no problem as demonstrated by the simulation study (Table 1) because the confidence sets are reliant on the value of the function $CD(\theta)$ instead of tail areas. In conducting the simulation study for assessing the coverage properties of the resulting confidence intervals, we considered $n=d=10$ and set $S=6$. In this example, the average lengths of the 0.95, 0.90 , 0.85 and 0.80 confidence intervals based on the LRT were 2.56, 2.88, 3.08 and 3.31, respectively, while those based on the CD-Box were 1.22, 1.37, 1.55 and 1.84, respectively.
@@ 4.4 Ricker’s Model
We conduct two experiments: first, we assume that only the log-growth rate is unknown and consider as summary statistics the median of counts and the quantiles of level 0.25, 0.75. For the second experiment, both the parameters $log(r)$ and $\sigma^{2}$ were considered unknown, and we used as the set of summary statistics the whole time series minus the first observation, thus of length $d=19$. The number of pseudo-samples for each proposals were $S=2$ and $S=10$ in the two experiments, respectively. The empirical coverages of the confidence sets are conformal with the nominal (Table 1). Two examples of confidence regions obtained for two independent draws from the model with parameters $log(r)=2$ and $\sigma^{2}=2$ are reported in Figure 5.